OpenStudy (anonymous):

CALCULUS PROBBBBBSSSS! f(x)= x^4 - 2x^3 - x^2 - 4x + 3 I= [0,4] MAX AND MIN

OpenStudy (anonymous):

find derivative set it equal to 0

OpenStudy (anonymous):

I'm so fast that last night I turned off the light switch in my hotel room and was in bed before the room was dark.

OpenStudy (anonymous):

also, since this is a closed interval you're looking at, evaluate at the enpoints of the interval...

OpenStudy (anonymous):

@satellite73 , Cassius Clay

OpenStudy (anonymous):

i got 4x^3 - 6x^2 -2x - 4 now what??

OpenStudy (anonymous):

aka imranmeah

OpenStudy (anonymous):

??? never heard of 'em

OpenStudy (anonymous):

try something obvious like x = 1, x = -1, x = 2 you will get it on the third try

OpenStudy (anonymous):

shut up satellite73 THIS IS MATH ONLY

OpenStudy (anonymous):

how you are supposed to know that i am not sure, but \(x=2\) is a root of the derivative

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

\[f'(x)=4x^3 - 6x^2 -2x - 4 \] \[f'(2)=0\] is the critical point

OpenStudy (anonymous):

so your last job is to take \[f(0), f(2), f(4)\] to find the max and min

OpenStudy (anonymous):

how did you get x = 2???

OpenStudy (anonymous):

Also, graphing it might help at the end points 0 and 4. See the attached

OpenStudy (anonymous):

\[4 x^3-6 x^2-2 x-4=2 (x-2) \left(2 x^2+x+1\right)\]

OpenStudy (anonymous):

i cheated. but you should try the obvious possibilities first. if you have a rational zero, then the numerator must divide 4 and the denominator must also divide 4. since this problem was cooked up by some math teacher 2 works out. that is why i suggested try 1 and -1 first because they are easiest to check. you can try to factor as eliassab did, but honestly i have no idea how to factor a cubic polynomial without knowing one of the roots.

OpenStudy (anonymous):

Knowing one root r than dividing by (x -r) will help you factor your cubic.

OpenStudy (anonymous):

THANK YALLLLL!!!! much love!!